# Revisiting the Canonicalization for Fast and Accurate Crystal Tensor Property Prediction

### v3

@@ A.10 Experimental details
Piezoelectric 4998 0.43 3.09 $\text{C}/\text{m}^{2}$
@@
Experimental Settings. The experiments were performed using a single NVIDIA GeForce RTX 3090 GPU. For benchmarking, we directly utilized the codebases for MEGNET, ETGNN, and GMTNet as provided by Yan et al. [51]. For each property, the dataset is split into training, validation, and test sets in an 8:1:1 ratio. During model training, Huber loss, AdamW [24], a weight decay of $10^{-5}$, and polynomial learning rate decay were employed.
Hyperparameter settings of GoeCTP. When constructing the crystal graph, we used the 16th nearest atom to determine the cutoff radius. For edge embeddings, we used an RBF kernel with $c=0.75$ and values ranging from --4 to 0, which maps $-c/||\mathbf{e}_{ij}||_{2}$ to a 512-dimensional vector. In the dielectric and piezoelectric tensor prediction task, the 512-dimensional vector is mapped to a 128-dimensional vector through a non-linear layer, while in the elastic tensor prediction, it is mapped to a 256-dimensional vector. For the prediction module, eComFormer, in the dielectric and piezoelectric tensor prediction task, we used 4 node-wise transformer layers and 1 node-wise equivariant updating layer, whereas in the elastic tensor prediction, we used 2 node-wise transformer layers and 1 node-wise equivariant updating layer. For both dielectric, piezoelectric, and elastic tensor tasks, the learning rate was set to 0.001, with 200 epochs and a batch size of 64. In the Reverse R&R module, for the dielectric tensor prediction task, the 128-dimensional vector features output by the prediction module are mapped to a 6-dimensional vector through two non-linear layers, then reconstructed into a $3\times 3$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ obtained from the R&R module to achieve $O(3)$ equivariant output. For the piezoelectric prediction task, the 18-dimensional vector features output by the prediction module are reconstructed into a $3\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output. For the elastic tensor prediction task, the 256-dimensional vector features output by the prediction module are mapped to a 36-dimensional vector through two non-linear layers, then reconstructed into a $6\times 6$ matrix and combined with the orthogonal matrix $\mathbf{Q}$ from the R&R module to achieve $O(3)$ equivariant output.
@@
Hyperparameter settings of GMTNet, ETGNN, and MEGNET. Following Yan et al. [51], we trained GMTNet, ETGNN, and MEGNET for 200 epochs using Huber loss with a learning rate of 0.001 and Adam optimizer with $10^{-5}$ weight decay across all tasks. The same polynomial learning rate decay scheduler is used in all experiments.
A.11 Additional results.
We assessed the performance of GoeCTP when combined with iComFormer [50] and CrystalFormer [42] (denoted as GoeCTP (iCom.) for the combination) on both the dielectric, piezoelectric, and elastic tensor datasets. As presented in Table 12, on the dielectric tensor dataset, although the prediction accuracy of GoeCTP (iCom.) and GoeCTP (Crys.) did not exceed that of GoeCTP (eCom.), GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.). As shown in Table 13, similar to the dielectric tensor dataset, GoeCTP (eCom.) achieved the best prediction accuracy on the piezoelectric dataset, but GoeCTP (iCom.) demonstrated slightly higher computational efficiency compared to GoeCTP (eCom.). In contrast, as shown in Table 14, on the elastic tensor dataset, GoeCTP (iCom.) and GoeCTP (Crys.) achieved superior prediction quality relative to GoeCTP (eCom.), albeit with lower efficiency.
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
Fnorm $\downarrow$ 3.71 3.40 3.28 3.23 3.40 3.53
EwT 25% $\uparrow$ 75.8% 82.6% 83.3% 83.2% 81.7% 80.1%
EwT 10% $\uparrow$ 38.9% 49.1% 56.0% 56.8% 53.8% 52.9%
EwT 5% $\uparrow$ 18.0% 25.3% 30.5% 35.5% 32.3% 30.6%
Total Time (s) $\downarrow$ 663 1325 1611 616 535 645
Time/batch (s) $\downarrow$ 0.052 0.104 0.126 0.048 0.042 0.202
@@
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
Fnorm $\downarrow$ 0.504 0.515 0.450 0.431 0.448 0.457
EwT 25% $\uparrow$ 39.3% 36.3% 45.5% 46.9% 45.5% 42.7%
EwT 10% $\uparrow$ 27.1% 20.5% 38.1% 42.9% 44.9% 42.1%
EwT 5% $\uparrow$ 11.4% 13.0% 34.3% 39.7% 38.3% 41.5%
Total Time (s) $\downarrow$ 943 1220 5771 938 909 2787
Time/batch (s) $\downarrow$ 0.074 0.095 0.45 0.073 0.071 0.218
Table 13: Additional comparison of performance metrics on piezoelectric dataset.
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.)
Fnorm $\downarrow$ 143.86 123.64 117.62 107.11 102.80 107.44
EwT 25% $\uparrow$ 23.6% 32.0% 36.0% 42.5% 46.7% 43.5%
EwT 10% $\uparrow$ 3.0% 3.8% 7.6% 15.3% 18.6% 15.8%
EwT 5% $\uparrow$ 0.5% 0.5% 2.0% 7.2% 8.2% 7.9%
Total Time (s) $\downarrow$ 2899 4448 $>$ 36000 2422 4035 7891
Time/batch (s) $\downarrow$ 0.226 0.348 $>$ 2.813 0.189 0.315 0.616
@@
A.12 The limitations of our method and its broader impact
Limitations. (1)As discussed in the main text, GoeCTP is a plug-and-play $O(3)$-equivariant framework designed to enhance the backbone network’s ability to achieve equivariant predictions. Consequently, the performance of our method is inherently dependent on the capabilities of the backbone network. If the backbone lacks sufficient predictive power, the combined framework may not fully surpass the state-of-the-art tensor prediction networks in all evaluation metrics. For instance, as shown in Table 12, when our framework is integrated with CrystalFormer [42], although the computational speed significantly exceeds that of GMTNet, the Fnorm, EwT 25%, and EwT 10% values are lower than those achieved by GMTNet.
(2)As demonstrated in Higham [12] and Proposition 3.5, polar decomposition applied to a $3\times 3$ invertible matrix $\mathbf{L}$ produces a unique $3\times 3$ orthogonal matrix $\mathbf{Q}$. For 3D crystal structures, the lattice matrix $\mathbf{L}$ is always full-rank (i.e. invertible), therefore, ensuring the applicability of our method to 3D crystal systems. However, for certain special cases, such as 2D crystals with single-layer structures [32, 39], the rank of the lattice matrix $\mathbf{L}$ may be less than 3. In these scenarios, directly applying polar decomposition may not yield a unique $3\times 3$ orthogonal matrix $\mathbf{Q}$ and a unique canonical form $\mathbf{H}$. This limitation could cause our method to fail, as crystals with different spatial orientations cannot be consistently adjusted to a unique canonical form, thereby preventing the method from achieving the desired $O(3)$-equivariance.
Broader impact. This work can be applied to screen existing crystal material databases for potential material candidates, facilitating the discovery of new materials with desirable tensor properties, such as high-dielectric materials. Therefore, our work has the potential to make a significant impact on materials science.
A.13 How to further utilize the tensor properties symmetry
@@
Utilizing symmetry constraints for zero elements. To demonstrate the role of symmetry constraints, we present an example of the GoeCTP prediction results in Table 15. Within a reasonable margin of error, the predictions align well with the expected constraints.
@@
$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.252&0.016&0.008\\ 0.016&2.230&0.007\\ 0.008&0.007&2.262\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
@@
For a dielectric tensor, we evaluate the prediction of zero elements by defining a threshold as 1% of the average value of non-zero elements in the labels. Predictions are deemed successful if they meet this threshold. The results are as shown in Table 16. It can be observed that our method predicts most zero elements accurately, though not flawlessly. Since our advantage lies in transferring equivariance through an external framework, there are no restrictions on the model itself. Therefore, to achieve a higher success rate for zero elements, we added a ReLU activation function to the output layer of the network to improve the success rate (this applies only to cases where tensor elements are greater than or equal to zero). The results after retraining GoeCTP are as shown in Table 17, Table 18, and Table 19. Our method predicts zero elements more accurately while maintaining overall performance.
@@
$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.237&0.000&0.000\\ 0.000&2.283&0.000\\ 0.000&0.000&2.228\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
@@
This simple modification enables GoeCTP to more accurately predict the zero elements in dielectric tensors caused by the space group constraints.
@@
In the first method, only the independent components of the tensor properties are predicted. For example, for a cubic crystal, we predict only one independent component and then use the mask from Table 8 to reconstruct it into a $3\times 3$ tensor). In Table 15, GoeCTP was modified to directly predict independent components. The new results are shown in Table 20, which fully satisfy the symmetry constraints. As shown in Table 21, the overall prediction performance indicates that directly predicting the independent components improves the model’s Fnorm performance on the dielectric dataset. Since this method directly predicts independent components, it is more suitable for cases like the dielectric tensor, where the independent components exhibit relatively simple structures. However, for more complex datasets, such as those involving piezoelectric and elastic tensors, the model’s performance using this method may degrade. In such scenarios, the second method provides a more viable alternative.
@@
$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.357&0.000&0.000\\ 0.000&2.357&0.000\\ 0.000&0.000&2.357\end{pmatrix}$ $\bm{\varepsilon}=\begin{pmatrix}\bm{\varepsilon}_{11}&0&0\\ 0&\bm{\varepsilon}_{11}&0\\ 0&0&\bm{\varepsilon}_{11}\end{pmatrix}$
@@
In the second method, once the network training is completed, the network output is weighted using the standard tensor property format, referred to as a mask, as detailed in Table 8. The application of this method to the piezoelectric and elastic datasets is demonstrated in Tables 22 and 23. The results indicate that incorporating symmetry-based weighting effectively improves the performance of GoeCTP.
GoeCTP GoeCTP (weighted)
Fnorm $\downarrow$ 0.431 0.389
EwT 25% $\uparrow$ 46.9% 48.5%
EwT 10% $\uparrow$ 42.9% 46.7%
EwT 5% $\uparrow$ 39.7% 46.1%
Table 22: Performance comparisons between GoeCTP and GoeCTP (weighted by the standard tensor property form) on the piezoelectric dataset.
@@
A.14 Analysis of discontinuity in canonicalization
Previous studies [6] have shown that canonicalization can often introduce discontinuities into the model. To assess the impact of such discontinuities on our method, we analyze our proposed method from both theoretical and empirical perspectives.
From theoretical perspective: The following proposition, presented in Appendix B of Dym et al. [6], characterizes the existence of a continuous canonicalization under orthogonal group actions:
Proposition A.3.
@@ A.14 Analysis of discontinuity in canonicalization
According to Proposition A.3, for a matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$, its continuous canonicalization norm is given by $y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}$.
In this work, the polar decomposition is applied to the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H},\mathbf{L}\in\mathbb{R}^{3\times 3}$, where $\mathbf{Q}$ is an orthogonal matrix, and $\mathbf{H}$ is the positive semi-definite Gram matrix (which serves as the canonicalization norm). Based on the properties of positive semi-definite matrices, our method is actually equivalent to the canonicalization approach described in Proposition A.3, i.e. $y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}=\left((\mathbf{Q}\mathbf{H})^{\top}\mathbf{Q}\mathbf{H}\right)^{1/2}=\left(\mathbf{H}^{\top}\mathbf{H}\right)^{1/2}=\mathbf{H}$. ∎
@@
From empirical perspective: To further demonstrate how this discontinuity affects the accuracy of tensor property predictions, we conducted a set of comparative experiments in which we replaced the polar decomposition in GoeCTP with QR decomposition [22] for evaluation. In this QR decomposition case, the lattice matrix is decomposed as $\mathbf{L}=\mathbf{Q}\mathbf{R}$, where $\mathbf{Q}$ is an orthogonal matrix, and $\mathbf{R}$ is an upper triangular matrix that serves as the canonicalization norm. However, $\mathbf{R}$ does not satisfy Proposition A.3, and therefore does not constitute a continuous canonicalization approach.
@@
A.15 How to design other canonical forms
@@
A.16 Impact of different canonical form on tensor property prediction
As discussed in Appendix A.15, there exist infinitely many possible canonical forms. To investigate the impact of different canonical forms on tensor property prediction in crystalline materials, we first present the results comparing the performance of GoeCTP with QR decomposition [22] and GoeCTP without canonicalization. The corresponding results are summarized in the table below:

### v4

@@ A.6 Experimental details
Piezoelectric 2701 0.79 4.03 $\text{C}/\text{m}^{2}$
@@
Experimental Settings. The experiments were performed using a single NVIDIA GeForce RTX 3090 GPU. For benchmarking, we directly utilized the codebases for ETGNN and GMTNet as provided by yanspace. For each property, the dataset is split into training, validation, and test sets in an 8:1:1 ratio (yanspace).
[…]
Hyperparameter settings of GoeCTP (C. Fra.). For all tensor prediction tasks, the configuration of the prediction module, CrystalFramer (ito2025rethinking), including the input atom-embedding layer, self-attention blocks, and other architectural components, was kept identical to the setup described in the original work (ito2025rethinking) on the JARVIS dataset. For the dielectric tensor prediction task, the model was trained using an L1 loss function with a learning rate of 0.002, a batch size of 256, and 200 training epochs, following a polynomial learning rate decay schedule. For the piezoelectric tensor prediction task, the model employed an L1 loss function with a learning rate of 0.002, a batch size of 512, and 1000 training epochs, using a warm-up-free inverse square root schedule (huang2020improving; ito2025rethinking). For the elastic tensor prediction task, the model adopted the Huber loss function with a learning rate of 0.001, a batch size of 64, and 500 training epochs, along with a polynomial learning rate decay schedule. Across all experiments, the AdamW optimizer (loshchilovdecoupled) was used with parameters $(\beta_{1},\beta_{2})=(0.9,0.98)$ and a weight decay of $10^{-5}$.
@@
Hyperparameter settings of GMTNet and ETGNN. Following yanspace, we trained GMTNet and ETGNN for 200 epochs using Huber loss with a learning rate of 0.001 and Adam optimizer with $10^{-5}$ weight decay across all tasks. The same polynomial learning rate decay scheduler is used in all experiments. For the dielectric tensor prediction task and the piezoelectric tensor prediction task, the network configuration of GMTNet followed that described in the original paper, comprising two layers for node-invariant feature updating and three layers for equivariant message passing. For the elastic tensor prediction task, the configuration was adjusted to include two layers for node-invariant feature updating and five layers for equivariant message passing.
A.7 Additional results.
We assessed the performance of GoeCTP when combined with iComFormer (yancomplete) and CrystalFormer (taniaicrystalformer) on both the dielectric and elastic tensor datasets. We additionally included the baseline method MEGNet (chen2019graph; morita2020modeling) in our comparative experiments. Additional comparison results for the dielectric tensor dataset are presented in Table 11, and those for the elastic tensor dataset are presented in Table 12.
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.) GoeCTP (C. Fra.)
Fnorm $\downarrow$ 3.71 3.40 3.28 3.23 3.40 3.53 3.05
EwT 25% $\uparrow$ 75.8% 82.6% 83.3% 83.2% 81.7% 80.1% 86.4%
EwT 10% $\uparrow$ 38.9% 49.1% 56.0% 56.8% 53.8% 52.9% 62.6%
EwT 5% $\uparrow$ 18.0% 25.3% 30.5% 35.5% 32.3% 30.6% 43.5%
Total Time (s) $\downarrow$ 663 1325 1611 616 535 645 1976
@@
MEGNET ETGNN GMTNet GoeCTP (eCom.) GoeCTP (iCom.) GoeCTP (Crys.) GoeCTP (C. Fra.)
Fnorm $\downarrow$ 143.86 123.64 117.62 107.11 102.80 107.44 95.98
EwT 25% $\uparrow$ 23.6% 32.0% 36.0% 42.5% 46.7% 43.5% 49.7%
EwT 10% $\uparrow$ 3.0% 3.8% 7.6% 15.3% 18.6% 15.8% 20.3%
EwT 5% $\uparrow$ 0.5% 0.5% 2.0% 7.2% 8.2% 7.9% 9.8%
Total Time (s) $\downarrow$ 2899 4448 >> 36000 2422 4035 7891 17197
@@
A.8 How to further utilize the tensor properties symmetry
@@
Utilizing symmetry constraints for zero elements. To demonstrate the role of symmetry constraints, we present an example of the GoeCTP (eCom.) prediction results in Table 13. Within a reasonable margin of error, the predictions align well with the expected constraints.
@@
$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.252&0.016&0.008\\ 0.016&2.230&0.007\\ 0.008&0.007&2.262\end{pmatrix}$ $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&0\\ 0&\boldsymbol{\varepsilon}_{11}&0\\ 0&0&\boldsymbol{\varepsilon}_{11}\end{pmatrix}$
@@
For a dielectric tensor, we evaluate the prediction of zero elements by defining a threshold as 1% of the average value of non-zero elements in the labels. Predictions are deemed successful if they meet this threshold. The results are as shown in Table 14. It can be observed that our method predicts most zero elements accurately, though not flawlessly. Since our advantage lies in transferring equivariance through an external framework, there are no restrictions on the model itself. Therefore, to achieve a higher success rate for zero elements, we added a ReLU activation function to the output layer of the network to improve the success rate (this applies only to cases where tensor elements are greater than or equal to zero). The results after retraining GoeCTP (eCom.) are as shown in Table 15, Table 16, and Table 17. Our method predicts zero elements more accurately while maintaining overall performance.
@@
$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.237&0.000&0.000\\ 0.000&2.283&0.000\\ 0.000&0.000&2.228\end{pmatrix}$ $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&0\\ 0&\boldsymbol{\varepsilon}_{11}&0\\ 0&0&\boldsymbol{\varepsilon}_{11}\end{pmatrix}$
@@
This simple modification enables GoeCTP (eCom.) to more accurately predict the zero elements in dielectric tensors caused by the space group constraints.
@@
In the first method, only the independent components of the tensor properties are predicted. For example, for a cubic crystal, we predict only one independent component and then use the mask from Table 7 to reconstruct it into a $3\times 3$ tensor). In Table 13, GoeCTP (eCom.) was modified to directly predict independent components. The new results are shown in Table 18, which fully satisfy the symmetry constraints. As shown in Table 19, the overall prediction performance indicates that directly predicting the independent components improves the model’s Fnorm performance on the dielectric dataset. Since this method directly predicts independent components, it is more suitable for cases like the dielectric tensor, where the independent components exhibit relatively simple structures. However, for more complex datasets, such as those involving piezoelectric and elastic tensors, the model’s performance using this method may degrade. In such scenarios, the second method provides a more viable alternative.
@@
$\begin{pmatrix}2.258&0&0\\ 0&2.258&0\\ 0&0&2.258\end{pmatrix}$ $\begin{pmatrix}2.357&0.000&0.000\\ 0.000&2.357&0.000\\ 0.000&0.000&2.357\end{pmatrix}$ $\boldsymbol{\varepsilon}=\begin{pmatrix}\boldsymbol{\varepsilon}_{11}&0&0\\ 0&\boldsymbol{\varepsilon}_{11}&0\\ 0&0&\boldsymbol{\varepsilon}_{11}\end{pmatrix}$
@@
In the second method, once the network training is completed, the network output is weighted using the standard tensor property format, referred to as a mask, as detailed in Table 7. The application of this method to the elastic datasets is demonstrated in Table 20. The results indicate that incorporating symmetry-based weighting effectively improves the performance of GoeCTP.
@@
A.9 Analysis of discontinuity in canonicalization
Previous studies (dymequivariant) have shown that canonicalization can often introduce discontinuities into the model. To assess the impact of such discontinuities on our method, we analyze our proposed method from both theoretical and empirical perspectives.
From theoretical perspective: The following proposition, presented in Appendix B of dymequivariant, characterizes the existence of a continuous canonicalization under orthogonal group actions:
Proposition 3.
@@ A.9 Analysis of discontinuity in canonicalization
According to Proposition 3, for a matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$, its continuous canonicalization norm is given by $y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}$.
In this work, the polar decomposition is applied to the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H},\mathbf{L}\in\mathbb{R}^{3\times 3}$, where $\mathbf{Q}$ is an orthogonal matrix, and $\mathbf{H}$ is the positive semi-definite Gram matrix (which serves as the canonicalization norm). Based on the properties of positive semi-definite matrices, our method is actually equivalent to the canonicalization approach described in Proposition 3, i.e. $y_{\mathbf{L}}=\left(\mathbf{L}^{\top}\mathbf{L}\right)^{1/2}=\left((\mathbf{Q}\mathbf{H})^{\top}\mathbf{Q}\mathbf{H}\right)^{1/2}=\left(\mathbf{H}^{\top}\mathbf{H}\right)^{1/2}=\mathbf{H}$. ∎
@@
From empirical perspective: To further demonstrate how this discontinuity affects the accuracy of tensor property predictions, we conducted a set of comparative experiments in which we replaced the polar decomposition in GoeCTP (eCom.) with QR decomposition (linequivariance) for evaluation. In this QR decomposition case, the lattice matrix is decomposed as $\mathbf{L}=\mathbf{Q}\mathbf{R}$, where $\mathbf{Q}$ is an orthogonal matrix, and $\mathbf{R}$ is an upper triangular matrix that serves as the canonicalization norm. However, $\mathbf{R}$ does not satisfy Proposition 3, and therefore does not constitute a continuous canonicalization approach.
@@
A.10 How to design other canonical forms
@@
A.11 Impact of different canonical form on tensor property prediction
As discussed in Appendix A.10, there exist infinitely many possible canonical forms. To investigate the impact of different canonical forms on tensor property prediction in crystalline materials, we first present the results comparing the performance of GoeCTP (eCom.) with QR decomposition (linequivariance) and GoeCTP (eCom.) without canonicalization. The corresponding results are summarized in the table below: